OpenStudy (anonymous):

Evaluate using the substitution method (integral sign) ln(9x)/x dx

OpenStudy (anonymous):

Let u = ln(9x) du = 1 / x integral of ln(9x) / x dx = integral of u du = 1/2u^2 Substituting back, we get 1/2 * (ln(9x))^2

OpenStudy (anonymous):

+C